OpenStudy (anonymous):

Find the slope of a line parallel to the line through the given points. P(-2, 4), Q(6, -2) @nikato I got 4/3. Is this right?

OpenStudy (anonymous):

I did \[\frac{ y2-y1 }{ x2-x1 } \] and got \[\frac{ -6 }{ 8 }\] simplified to \[\frac{ -3 }{ 4}\] parallel would be \[\frac{ 4 }{ 3 }\] right?

OpenStudy (nikato):

No, if 2 lines are parallel, they would have the same slope

OpenStudy (nikato):

S it'll be just -3/4

OpenStudy (nikato):

4/3 would be the slope of the line perpendicular to it

OpenStudy (anonymous):

but.... parallel is being an equal distance apart everywhere

OpenStudy (nikato):

But ur just finding the slope, right?

OpenStudy (anonymous):

yes

OpenStudy (nikato):

Yes, and u found that it was -3/4 and the line parallel to that line would have the same slope, so that line also has a slope of -3/4

OpenStudy (anonymous):

Oh, okay. Thank you!

OpenStudy (nikato):

Yea, no problem. Are u sure u understand it now?

OpenStudy (anonymous):

I think I'm just a bit confused on what the slope actually is :T

OpenStudy (anonymous):

I mean parallel means they don't touch, I don't understand how they could have the same anything

OpenStudy (nikato):

Slope is kinda like how much the line slants, like its steepness

OpenStudy (nikato):

Yes, parallel means they never touch. And in order for that to happen, they must have the same "steepness"

OpenStudy (anonymous):

Oh, now I see. But perpendicular would have opposite slopes, right?

OpenStudy (anonymous):

reciprocal with opposite sign would be for perpendicular

OpenStudy (anonymous):

...what?

OpenStudy (anonymous):

It would be the opposite slope right? e.g. -1/3 and 1/3

OpenStudy (anonymous):

oops I meant -3/-1 and 1/3

OpenStudy (anonymous):